# Interpretable Company Similarity with Sparse Autoencoders

### v1

@@ 2.6 Pairs Trading
The dataset is divided into an in-sample period (Jan 2002–Dec 2013) and an out-of-sample period (Jan 2014–Dec 2020), with clusters $G_{k}$ ($G_{\text{C-T1}}$, SIC, BISC, and their joined variations) identified using cluster models based on 2011, 2012, and 2013. The pairs trading strategy involves the following steps:
@@
    Pre-selection: For each cluster $C_{i}\in G_{k}$, stock pairs are filtered if the Pearson correlation of their monthly logged returns exceeds 0.9 during the in-sample period. This step ensures that only pairs with strong linear relationships are considered.
@@
    Cointegration Testing: Using historical adjusted close data (Jan 2002–Dec 2013) for the pre-selected pairs while adjusting for corporate action, an Engle-Granger cointegration test is conducted with the Augmented Dickey-Fuller (ADF) statistic to determine whether the residual spread is stationary. Pairs with a p-value below 0.05 are considered co-integrated.
@@
    Ranking and Selection: Co-integrated pairs are ranked by their ADF test p-values [^7] We choose pairs with lower p-values, as they indicate stronger cointegration relationships, with the top 100 pairs [^8] When the list reaches 200 pairs, it will stop searching to conserve computing time, and perform ranking retained for out-of-sample evaluation.
In the out-of-sample evaluation (Jan 2014–Dec 2020), we simulated live trading of co-integrated pairs using a mean-reversion strategy. Following Miao (2014), we assumed zero transaction costs, opening positions when the residual spread deviated beyond ±$\pm 1\sigma$ its mean, and closing when the spread reverted to the mean. A stop-loss mechanism is triggered if the spread exceeds ±$\pm 2\sigma$. We allocated fixed capital per pair, normalizing PnL to ensure comparability across clustering methods.
@@ 3 Results
3.1 Classifier Results
Logistic Regression and the SNN are trained on data from $year_{i}$ and tested on data from $year_{i+1}$, with the goal of assigning probabilities to each company pair. The average correlation values between the generated similarity scores and the return correlations over 24 years are presented below.
Logistic Classifier (LC): The Logistic Classifier with a custom loss function significantly outperforms the SIC benchmark in predicting financial correlations, achieving a correlation of 0.081 compared to SIC’s correlation of -0.006. This suggests that custom loss functions tailored to financial contexts can enhance predictive accuracy.
Siamese Neural Network (SNN): The Siamese Neural Network outperforms the logistic regressor by effectively capturing non-linear relationships within our features. The training process exhibits consistent convergence, avoiding premature loss plateaus. Despite the simplicity and low dimensionality of the features, this out-performance suggests that the approach is both scalable and robust for capturing complex financial similarities. Additionally, the Pearson correlation coefficient between the predicted similarity scores and the actual returns correlation demonstrates a relatively strong positive relationship and validates the effectiveness of the network (Table 1).
Classifier Algorithm Average Correlation
Siamese Neural Network 0.179
Logistic Classifier 0.081
Table 1: Average Pairwise Correlations of the Classifier Models (1996 to 2020). For reference, when trying to predict the correlation with the Logistic Classifier using SIC the correlation is -0.006, and 0.059 with BISC. We do not compute it for the Siamese compression network since SIC codes can’t be compressed.
@@ 3.2 Clustering Results
Figure 2: Mean correlation of C-TS (Siamese) Cluster vs Benchmarks between 1996-2020. Joint model significantly outperforming benchmarks, which may suggest sparse features and SIC capture complementary, non-overlapping information.
We compare PnL of the clusters we obtain using sparse features against SIC codes and combined benchmarks, drawn from intersections (Figure 3). The results demonstrate that pairs derived from our Sparse Features outperform SIC and BISC in average pairwise correlation. Furthermore, the intersection of clusters with SIC codes leads to the most granular characterization and strongest performance. This highlights the complementary nature of traditional and feature-based classification systems. Further, we check the mean pairwise correlation and Sharpe ratio within each cluster (Table 2).
Clustering Algorithm Avg Corr. Sharpe
C-T1 0.29 -10.3
C-T3 0.27 10.74
C-TA 0.24 12.54
C-TS 0.25 12.97
C-TK 0.33 -
C-T1 $\cap$ SIC 0.45 -
C-T1 $\cap$ BISC 0.37 -
C-T3 $\cap$ SIC 0.39 14.41
C-T3 $\cap$ BISC 0.35 14.41
C-TA $\cap$ SIC 0.41 -
C-TA $\cap$ BISC 0.39 -
C-TS $\cap$ SIC 0.43 -
C-TS $\cap$ BISC 0.32 13.88
C-TK $\cap$ SIC 0.39 -
C-TK $\cap$ BISC 0.36 -
C-LC1_SIC 0.15 -
C-LC1_BISC 0.20 -
SIC 0.23 12.76
BISC 0.19 13.04
Population 0.16 - -
Table 2: Performance comparison of clustering algorithms across 1996-2020. Avg Corr – average pairwise correlation.
@@ 3.3 Pairs Trading Results
During the live trading simulation, cumulative PnL (Smith and Doe, 2021) was recorded for evaluation. Results are reported in Figure 3.
Figure 3: Cumulative PnL Trajectory Comparison between C-TS Cluster vs Benchmarks: In particular, notice the intersection between C-TC and BISC (major group classification) significantly outperforms benchmarks.
While the co-integration test ensures statistical validity, the strategy’s profitability relies on the underlying structural fundamental similarities between companies such as business models and operational structures (Hong and Hwang, 2021). The aforementioned study establishes that portfolios of pairs with higher similarities in firm characteristics minimize spread non-convergence risk, outperforming fundamentally less similar pairs. Our clustering approach outperforms SIC and BISC groups in PnL and risk-adjusted metrics, suggesting our method captures more fundamental aspects of company similarity.
@@ 3.4 Limitations
The following are the main limitations: Firstly, while the distribution of the similarity metric is quasi-normal, it exhibits high kurtosis, making it more difficult to assign a stock pair to a meaningful cluster. Our analysis primarily concerns the tails of the distribution, which contain the most similar stock pairs. This limits the number of clusters we are able to analyze and diminishes the scope of the model’s applicability. Further, the exclusion of delisted stocks introduces survivorship bias into the analysis, reducing its ability to accurately represent real-world market dynamics. Secondly, we rely on a learned similarity function that is based on correlations. This is not a metric in the strict sense. While this approach has precedent (Mantegna, 1999), we are thinking about alternatives such as (Mantegna, 2015). Thirdly, when trading co-integrated pairs, we select tickers purely based on quantitative measures, but one could argue that a strategy should be built with pairs selected via a stronger economic rationale.
@@ 4 Conclusions
We find that using SAE features is an effective and interpretable method for computing company similarity and explaining returns. Stress-testing, portfolio diversification, and other hedging-related strategies could all be subjects of further investigation using our clustering approach. We remain excited for future work in this direction, and will release all our datasets and code in the camera ready.

### v2

@@ 2.6 Pairs Trading
The dataset is divided into an in-sample period (Jan 2002–Dec 2013) and an out-of-sample period (Jan 2014–Dec 2020), with clusters $G_{k}$ ($G_{\text{C-T1}}$, SIC-codes, BISC, and their joined variations). The pairs trading strategy involves the following steps:
@@
    Pre-selection: For each cluster $C_{i}\in G_{k}$, stock pairs are filtered if the Pearson correlation of their monthly logged returns exceeds 0.9 during the in-sample period.
@@
    Co-integration Testing: Using historical adjusted close data (Jan 2002–Dec 2013) for the pre-selected pairs while standardized for corporate action, an Engle-Granger co-integration test is conducted with the Augmented Dickey-Fuller (ADF) statistic to determine whether the residual spread is stationary. Pairs with a p-value below 0.05 are considered co-integrated.
@@
    Ranking and Selection: Co-integrated pairs are ranked by their ADF test p-values [^9] We choose pairs with lower p-values, as they indicate stronger stationarity of the spread, reflecting a stronger co-integration relationship, with the top 80 pairs retained for out-of-sample evaluation (see Table 1). The effectiveness of co-integration within each cluster group $Gk$ is assessed by evaluating the Sharpe ratio of the entire portfolio. The Sharpe ratio quantifies risk-adjusted returns, measuring excess return per unit of risk. We prioritized this over absolute PnL, as higher Sharpe ratios enable leverage to amplify returns, provided practical constraints such as transaction costs and risk limits are effectively managed (Guasoni and Mayerhofer, 2018; Peters, 2011).
See Appendix D for Trading Logic Detail.
@@ 3.1 Clustering Results
Figure 2: Mean correlation of C-TS (Siamese) Cluster vs Benchmarks between 1996-2020. Joint model significantly outperforming benchmarks, suggesting sparse features and SIC-codes capture complementary, non-overlapping information.
[…]
Clustering Group $\mathbf{(G_{k})}$ $\text{MeanCorr}\mathbf{(G_{k})}$ (1996-2020) Sharpe Ratio
$\mathbf{G_{\text{C-TS}}}$ $\mathbf{0.246}$ $\mathbf{12.97}$
$\mathbf{G_{\text{C-TM}}}$ $\mathbf{0.266}$ $\mathbf{15.84}$
$G_{\text{BERT}}$ 0.198 12.68
$G_{\text{SBERT}}$ 0.219 9.34
$G_{\text{PaLM-gecko}}$ 0.219 9.97
$\mathbf{G_{\text{C-TS $\cap$ SIC }}}$ $\mathbf{0.454}$mathbf{-}-
$\mathbf{G_{\text{C-TS $\cap$ BISC}}}$ $\mathbf{0.327}$ $\mathbf{13.88}$
$\mathbf{G_{\text{C-TM $\cap$ SIC}}}$ $\mathbf{0.416}$mathbf{-}-
$\mathbf{G_{\text{C-TM $\cap$ BISC}}}$ $\mathbf{0.412}$mathbf{-}-
$G_{\text{BERT $\cap$ SIC}}$mathbf{-}- 12.11
[…]
$G_{\text{SBERT $\cap$ BISC}}$ 0.292 14.66
[…]
SIC 0.231 10.73
BISC 0.187 13.85
[…]
@@ 3.2 Pairs Trading Results
During the live trading simulation, cumulative PnL (profits) (Smith and Doe, 2021) and Sharpe ratios (risk-adjusted profits) were recorded for evaluation. Results are reported in Figure 3.
Figure 3: Cumulative PnL and Sharpe ratios comparison between $G_{\text{C-TM}}$ Cluster and Benchmarks: $G_{\text{C-TM}}$ significantly outperforms in terms of Sharpe
While the co-integration test ensures statistical validity, the strategy’s profitability relies on the underlying structural fundamental similarities between companies. Our clustering approach outperforms SIC & BISC groups in Sharpe ratio, suggesting our method captures more fundamental aspects of company similarity (Hong and Hwang, 2021).
3.3 Interpretability
Sparse features are understood to be interpretable (Huben et al., 2024), and our core contribution is showing that it is possible to match or improve upon existing methodologies (U.S. Occupational Safety and Health Administration, 2001; Vamvourellis et al., 2023) while imposing an interpretability constraint.
In practice, we use interpretability to ascertain the economic rationale underlying the most co-integrated stock pairs within each cluster $C_{i}$ in $G_{\text{C-TM}}$. This allows us to go beyond statistical tests on correlation and spread-stationarity, ensuring that the co-integration is supported by valid economic relationships (Hong and Hwang, 2021).
For each pair of stocks in our portfolio we first obtain an interaction feature vector $f_{i,j}$ by pairwise multiplying the feature vectors $s_{i},s_{j}$ representing the two company descriptions. Then we inspect the features that activate the most across all interaction vectors in our portfolio, and analyze them on the company descriptions of our most co-integrated pair using sae-vis (McDougall, 2024).
The co-integrated pair with the highest Sharpe ratio in $G_{\text{C-TM}}$ is UFP Technologies and Heico. Both share operational similarities in highly specialized manufacturing: UFP in custom-engineered products for medical and industrial applications; Heico produces aerospace, defense, and industrial components. Many interpretable and relevant features emerge, such as a ’renting’ feature 20155 highlighting both companies’ niche operations with logistical dependencies, including renting specialized facilities, exposing them to real estate fluctuations. We interpret the feature by examining its impact on the LLM’s logic in Appendix F.
@@ 3.4 Limitations
The following are the main limitations:
  1. 1.
    We do not fine-tune embedders, LLMs, nor SAEs. This is due to the high computational cost of supervised fine-tuning exceeding our computing budget, though fine-tuning could improve the performance of our method. This is an exciting direction for future work.
  2. 2.
    The exclusion of delisted stocks introduces survivorship bias into the analysis. This can be addressed with more complete data.
  3. 3.
    Our interpretability study is a proof of concept, limited by the availability of human feature annotators (there are tens of thousands of active features), and while we rely on the literature as support for the claim that SAE features are interpretable (Huben et al., 2024), our study could be expanded in the future.
@@ 4 Conclusions
We find that using SAE features is an effective and interpretable method for computing company similarity and explaining returns. Stress-testing, portfolio diversification, and other hedging-related strategies could all be subjects of further investigation using our clustering approach. We will release all our datasets and code in the camera ready.
Appendix A Clustering Embeddings
For BERT, we used bert-base-uncased from the transformers library. For SBERT, we used all-MiniLM-L6-v2 from the sentence_transformers library. For PaLM-gecko, we used textembedding-gecko@003 from the vertexai library.
Chunking: In our methodology, for both BERT and SBERT, we followed Vamvourellis et al. (2023) and implemented a chunking mechanism to accommodate the models’ maximum token limit of 512. Specifically, company descriptions exceeding this limit were split into overlapping chunks of 512 tokens. The [CLS] embeddings of these chunks were averaged to generate a single document embedding of 1536 tokens. For PaLM-Gecko, we leveraged its extended context window of 3072 tokens and directly processed the descriptions without chunking.
The pipeline below is optimised through Optuna’s Tree-structured Parzen Estimator (TPE) sampler for Bayesian hyperparameter optimization. The objective function maximizes $\text{MeanCorr}(G_{k})$. This search is constrained to 150 trials and a maximum timeout of 9 hours to balance thoroughness and resource usage:
(a) Dimensionality Reduction with UMAP: Given the high dimensionality of the input embeddings (768-dimensional vectors derived from a BERT model), we employ Uniform Manifold Approximation and Projection (UMAP) umap to reduce these high-dimensional textual embeddings to a lower-dimensional space, preserving both local and global data structures. We optimize three UMAP parameters to improve the quality of the downstream clustering: (1) n_components (target dimensionality); (2) n_neighbors; and (3) min_dist. All embeddings are standardized and casted to float32 to ensure computational efficiency.
(b) Clustering with Spectral Clustering: After reducing dimensionality, we perform clustering using Spectral Clustering, which is capable of handling noise and complex cluster shapes, following Vamvourellis et al. (2023). We first construct an affinity matrix from a k-nearest neighbors (KNN) graph of the UMAP outputs. Spectral Clustering then operates on this graph’s eigenstructure to form clusters. The number of clusters (n_clusters) is tuned via Optuna, while the neighborhood size ($k$) is set to a constant of 5, following Vamvourellis et al. (2023).
(c) Temporal Cross-Validation: To evaluate the stability and temporal generalization of the resulting clusters, we employ temporal cross-validation. The dataset is split into chronological folds. This setup reduces temporal bias and assesses whether the identified cluster structure remains consistent over time. We used parallel processing to evaluate each fold.
Embedder Cluster Group $(G_{\text{embedder}})$ UMAP $n_{\text{components}}$ UMAP $n_{\text{neighbors}}$ UMAP min_dist
$G_{\text{BERT}}$ 7 119 0.109
$G_{\text{SBERT}}$ 7 79 0.012
$G_{\text{PaLM-gecko}}$ 6 40 0.120
Table 2: Optimized UMAP Thresholds for Embedders
Embedder Cluster Group $(G_{\text{embedder}})$ Spectral $n_{\text{clusters}}$ Spectral $n_{\text{neighbors}}$
$G_{\text{BERT}}$ 10 5
$G_{\text{SBERT}}$ 49 5
$G_{\text{PaLM-gecko}}$ 27 5
Table 3: Optimized Spectral Clustering Thresholds for Embedders
Appendix B Clustering Sparse Features
We optimize the cutoffs using Optuna in cross validation. We observe that nearby nodes in the Minimum Spanning Tree (MST) are close in value, indicating a smooth distribution of distances. This is evident in the plot of the edge-weight cutoff threshold (x-axis) against $\text{MeanCorr}(G_{k})$ of the resulting cluster groups (y-axis), which shows a continuous and stable trend. The smoothness of the distribution suggests that the underlying sparse features are meaningfully grouped, and the cutoff thresholds effectively preserve local structure while removing noise. This property supports the reliability of the generated clusters and aligns with the ultrametric framework, ensuring interpretability and robustness.
Sparse Feature Cluster Group $(G_{\text{sparse\_feature}})$ MST edge-weight cutoff
$G_{\text{C-TS}}$ 0.1940
$G_{\text{C-TM}}$ 0.2315
Table 4: Optimized MST edge-weight cutoff Thresholds for Sparse Feature Cluster Group
Appendix C Proof of Manhattan metric
To show that $M_{i,j}$ satisfies positive definiteness, we verify the following:
1. Non-negativity: By definition,
$M_{i,j}=\sum_{k=1}^{4000}\left|\text{PCA}_{k}(f_{i})-\text{PCA}_{k}(f_{j})\right|\geq 0,$
as absolute differences are always non-negative.
Zero only for identical points: If $M_{i,j}=0$, then each term in the sum must be zero, i.e.,
$\text{PCA}_{k}(f_{i})=\text{PCA}_{k}(f_{j})\quad\forall k.$
Since the PCA transformation is invertible for features in its span, this implies $f_{i}=f_{j}$. Conversely, if $f_{i}=f_{j}$, then all differences are zero, so $M_{i,j}=0$.
2. Symmetry: Using the property of absolute values,
$\left|\text{PCA}_{k}(f_{i})-\text{PCA}_{k}(f_{j})\right|=\left|\text{PCA}_{k}(f_{j})-\text{PCA}_{k}(f_{i})\right|\quad\forall k.$
Therefore,
$M_{i,j}=M_{j,i}.$
3. Triangle Inequality:
For each dimension $k$, for all $l$, the absolute value satisfies the triangle inequality:
$\left|\text{PCA}_{k}(f_{i})-\text{PCA}_{k}(f_{j})\right|$ $\leq\left|\text{PCA}_{k}(f_{i})-\text{PCA}_{k}(f_{l})\right|$
$\quad+\left|\text{PCA}_{k}(f_{l})-\text{PCA}_{k}(f_{j})\right|.$
Summing over all dimensions $k$ gives:
$\begin{array}[]{l}\displaystyle\sum_{k=1}^{4000}\left|\text{PCA}_{k}(f_{i})-\text{PCA}_{k}(f_{j})\right|\leq\sum_{k=1}^{4000}\left|\text{PCA}_{k}(f_{i})-\text{PCA}_{k}(f_{l})\right|\\ \displaystyle\quad+\sum_{k=1}^{4000}\left|\text{PCA}_{k}(f_{l})-\text{PCA}_{k}(f_{j})\right|.\end{array}$
This simplifies to:
$M_{i,j}\leq M_{i,l}+M_{l,j}.$
Thus, $M_{i,j}$ satisfies the triangle inequality.
Appendix D Trading details
In the out-of-sample evaluation (Jan 2014–Dec 2020), we simulated live trading of co-integrated pairs using a mean-reversion strategy.
Following Miao (2014), we assumed zero transaction costs, opening positions when the residual spread deviated beyond ±$\pm 1\sigma$ its mean, and closing when the spread reverted to the mean. A stop-loss mechanism is triggered if the spread exceeds ±$\pm 2\sigma$.
We allocated a fixed capital per pair and normalized the daily Total Portfolio Value (cash + unrealized PnL) to enable fair comparisons across clustering methods, regardless of the number of pairs. This normalization ensures that clusters with varying pair counts (e.g., 5 pairs vs. 60 pairs) have comparable portfolio value trajectories, where each pair is traded with an equivalent unit of capital (i.e., \$1). The daily normalized portfolio value series was subsequently used for Sharpe ratio calculations.
Appendix E Siamese Hyperparameters
Siamese Network was trained for 10 epochs with a learning rate of 0.001. The loss function used is Mean Squared Error (MSE). The hidden dimensions are 128 and 64 with a ReLu activation function.
Appendix F Interpretability of Feature 20155
Token Logit Contribution
tenant 0.28
tenants 0.24
Tenant 0.23
space 0.23
tenant 0.21
rents 0.21
shell 0.21
office 0.21
Table 5: Logit contribution on the top 8 tokens.